# Reassessing the gallium anomaly using exact electron wave functions

### v2

@@ II Determination of the neutrino capture cross section
where $\xi\left(5/2^{-}\right)$ and $\xi\left(3/2^{-}\right)$ are the phase-space coefficients, which take the values $\xi\left(5/2^{-}\right)=0.669$ and $\xi\left(3/2^{-}\right)=0.220$ for the ${}^{51}\textrm{Cr}$ neutrino source and $\xi\left(5/2^{-}\right)=0.696$ and $\xi\left(3/2^{-}\right)=0.264$ for the ${}^{37}\textrm{Ar}$ one [27]. The Gamow-Teller strength values for the excited contributions were obtained, following Ref. [27], from the experimental measurement of forward-angle (p,n) scattering [67] and from (${}^{3}\text{He},^{3}\text{H}$) scattering [68], leading to different estimations of about 6 % and 9 %, respectively. In both cases, we report the results in Table 3 for 51Cr and 37Ar neutrino sources.
@@ III Reavaluation of the anomaly
Our theoretical estimation of the number of expected IBD events, $N_{\mathrm{cal}}$, can be compared with the experimental measurements from GALLEX [3, 4, 5, 6], SAGE [7, 8] and BEST [24, 25], $N_{\mathrm{exp}}$, leading to a ratio $R=N_{\mathrm{exp}}/N_{\mathrm{cal}}$. We define the following $\chi^{2}$ function
@@
$\begin{split}\Delta\eta&=\left\{\begin{array}[]{cr}\Delta\eta_{+}(\mathrm{p,n})=0.047&\text{for}\,\eta>1,\\ \Delta\eta_{-}(\mathrm{p,n})=0.014&\text{for}\,\eta<1,\end{array}\right.\end{split}$ (26)
and $\Delta\eta(^{3}\mathrm{He},^{3}\mathrm{H})=0.029$ for the (${}^{3}\text{He},^{3}\text{H}$) one. The covariance matrix $V$ is obtained from the uncertainties on the $R_{k}$ ratios. The resulting best fit value for all the experiments is found to be $\overline{R}=0.814^{+0.029}_{-0.045}$, when considering the (p,n) excited contribution, see Fig. 2 (upper panel). Taking into account the non-parabolic behavior of the $\chi^{2}$, shown in Fig. 3, we obtain a $6.1\sigma$ anomaly. With the (${}^{3}\text{He},^{3}\text{H}$) excited contribution we obtain $\overline{R}=0.797^{+0.035}_{-0.034}$, also shown in Fig. 2 (lower panel), and a smaller $5.3\sigma$ anomaly. These results have been obtained by improving the calculation method used in Ref. [31] by removing the model uncertainties from the BEST data. These uncertainties are correlated among the experiments and are taken into account in the $\chi^{2}$ with the $\eta$ nuisance parameter. We also took into account the correlation of the systematic uncertainties of the BEST data. Moreover, to provide a conservative estimate of the anomaly, we reevaluate it considering the ground-state cross section in Eq. (17), whose best fit of the data is found to be $\overline{R}=0.860\pm 0.030$, which corresponds to a $4.5\sigma$ anomaly.
@@
Our result is in good agreement with previous ones [27, 28, 26], although the differences in the procedure. In fact, the averaging procedure compensates for slight differences in the electron wave function values adopted. It is worth stressing that in this work we use the same software and underlying assumptions when evaluating the exact electron wave functions in both the Fermi function and $\overline{f}_{\text{EC}}$ calculations in Eq. (17). Entering as a ratio, any systematic shift should cancel in the cross section determination, making the result more robust and self-consistent.
@@ IV Sterile Neutrinos
The simplest and most considered explanation of the gallium anomaly is the hypothesis of short-baseline active-sterile neutrino oscillations (see the recent review in Ref. [1]). Figure 4 shows the allowed regions in the space of the effective parameters of 3+1 short-baseline active-sterile oscillations (see Ref. [14]). Our results confirm those obtained in previous works [24, 71, 72, 73, 31, 11, 27, 20], which show that the gallium data, including the recent and most precise BEST data, can be explained by short-baseline active-sterile neutrino oscillations with a large mixing ($0.2\lesssim\sin^{2}2\vartheta_{ee}\lesssim 0.6$ at $2\sigma$, for $\Delta{m}^{2}_{41}\gtrsim 1\,\text{eV}^{2}$). Such large mixing is in tension with the bounds obtained from the analysis of reactor antineutrino data [74, 11, 75, 76, 77], the solar neutrino bound [78, 69], and the KATRIN neutrino mass measurement [70]. The solar and KATRIN bounds are shown in Fig. 4 for comparison. The reactor bound is more complicated, because it depends on the reactor antineutrino flux model and on the choice of the data set [74, 11]. These tensions disfavour the sterile neutrino explanation of the gallium anomaly, which is however still considered because so far nobody has found a plausible alternative explanation. [^11] Exotic explanations have been proposed in Refs. [79, 80, 81, 23, 30]. Some have been shown to be in tension with the reactor rate data [20].
@@ V Conclusions
In this work, we critically reassess the status of the gallium anomaly. By employing exact Dirac-Coulomb wave functions for both the bound and continuum electron states, a new averaging procedure and a refined statistical analysis, we reevaluate the neutrino capture cross-section on 71Ga, showing that a $6.1\sigma$ discrepancy is now achieved with the (p,n) excited state contribution, and $5.3\sigma$ with the (${}^{3}\text{He},^{3}\text{H}$) excited state contribution. Such a result is obtained by using the same software and underlying assumptions when evaluating the exact electron wave functions in both the Fermi function and electron density at the nucleus for the EC process, which entering as a ratio in the cross section avoid possible systematic shifts, making the result more robust and self-consistent. We also present the updated allowed region in the space of the effective parameters of 3+1 active-sterile neutrino mixing. We confirm the need for the explanation of the gallium data of a large active-sterile mixing which is in tension with the results of reactor antineutrino experiments [74, 11, 75, 76, 77], with the solar neutrino bound [78, 69], and the KATRIN neutrino mass measurement [70]. Hence, the solution of the gallium anomaly remains an open problem which may be solved by a new source experiment with a different detector [32, 82, 83].

### v3

@@ II Determination of the neutrino capture cross section
where $\xi\left(5/2^{-}\right)$ and $\xi\left(3/2^{-}\right)$ are the phase-space coefficients, which take the values $\xi\left(5/2^{-}\right)=0.669$ and $\xi\left(3/2^{-}\right)=0.220$ for the $^{51}\textrm{Cr}$ neutrino source and $\xi\left(5/2^{-}\right)=0.696$ and $\xi\left(3/2^{-}\right)=0.264$ for the $^{37}\textrm{Ar}$ one [27]. The Gamow-Teller strength values for the excited contributions were obtained, following Ref. [27], from the experimental measurement of forward-angle (p,n) scattering [70] and from ($^{3}\text{He},^{3}\text{H}$) scattering [71], leading to different estimations of about 6 % and 9 %, respectively. In both cases, we report the results in Table 3 for 51Cr and 37Ar neutrino sources. Such values can be compared to those reported by the most recent calculation available in the literature [27]. The ground-state cross sections agree within 0.8%, while, after applying the averaging procedure, our total cross sections result to be approximately 2% smaller than those reported in Ref. [27].
@@ III Reavaluation of the anomaly
Our calculation of the IBD cross section can be used to determine the number of expected IBD events, $N_{\mathrm{cal}}$, by considering the neutrino source activity, the atomic density of the target, the average neutrino path length through the absorption material and the exposure [72]. This value can be compared with the experimental measurements from GALLEX [3, 4, 5, 6], SAGE [7, 8] and BEST [24, 25], $N_{\mathrm{exp}}$, leading to a ratio $R=N_{\mathrm{exp}}/N_{\mathrm{cal}}$. Alternatively, one can retrieve directly the ratios $R$ for each experiment, which have been determined using a reference cross section [26], by rescaling them to correct for the difference between our cross section with the reference one. The two procedures lead to the same results, which are reported in Tab. 4. We define the following $\chi^{2}$ function
@@
[…]
and $\Delta\eta(^{3}\mathrm{He},^{3}\mathrm{H})=0.031$ for the ($^{3}\text{He},^{3}\text{H}$) one.
$R_{\rm gs}$ $R_{\rm(p,n)}$ $R_{\rm(^{3}\text{He},^{3}\text{H})}$
GALLEX Cr1 $1.049\pm 0.121$ $0.994\pm 0.115$ $0.966\pm 0.112$
GALLEX Cr2 $0.894\pm 0.121$ $0.847\pm 0.115$ $0.823\pm 0.112$
SAGE Cr $1.045\pm 0.132$ $0.991\pm 0.125$ $0.963\pm 0.122$
SAGE Ar $0.875\pm 0.105$ $0.824\pm 0.099$ $0.798\pm 0.096$
BEST Cr-R1 $0.870\pm 0.032(\text{unc})\pm 0.023(\text{cor})$ $0.825\pm 0.031(\text{unc})\pm 0.022(\text{cor})$ $0.802\pm 0.030(\text{unc})\pm 0.021(\text{cor})$
BEST Cr-R2 $0.843\pm 0.031(\text{unc})\pm 0.027(\text{cor})$ $0.799\pm 0.029(\text{unc})\pm 0.025(\text{cor})$ $0.777\pm 0.029(\text{unc})\pm 0.025(\text{cor})$
Table 4: Ratio between the experimental number of IBD events and the predicted one for each experimental measurement considered, together with the corresponding uncertainties, divided, when necessary, between correlated (cor) and uncorrelated (unc).
The covariance matrix $V$ is obtained from the uncertainties on the $R_{k}$ ratios. The resulting best fit value for all the experiments is found to be $\overline{R}=0.835^{+0.030}_{-0.048}$, when considering the (p,n) excited contribution, see Fig. 3 (upper panel). Taking into account the non-parabolic behavior of the $\chi^{2}$, shown in Fig. 4, we obtain a $5.5\sigma$ anomaly. With the ($^{3}\text{He},^{3}\text{H}$) excited contribution we obtain $\overline{R}=0.811^{+0.037}_{-0.035}$, also shown in Fig. 3 (lower panel), and a smaller $4.7\sigma$ anomaly. These results have been obtained by improving the calculation method used in Ref. [31] by removing the model uncertainties from the BEST data, to avoid double counting. These uncertainties are correlated among the experiments and are taken into account in the $\chi^{2}$ with the $\eta$ nuisance parameter. We also took into account the correlation of the systematic uncertainties of the BEST data. Moreover, to provide a conservative estimate of the anomaly, we reevaluate it considering the ground-state cross section in Eq. (17), whose best fit of the data is found to be $\overline{R}=0.881\pm 0.031$, which corresponds to a $3.8\sigma$ anomaly.
@@
Our result is in good agreement with previous ones [27, 28, 26], although differences in the procedure. In fact, the averaging procedure, which leads to a lower cross section value, compensates for the refined statistical analyses, leading to a similar significance of the anomaly. It is worth stressing that in this work we use the same software and underlying assumptions when evaluating the electron wave functions in both the Fermi function and $\overline{f}_{\text{EC}}$ calculations in Eq. (17). Entering as a ratio, any systematic shift should cancel in the cross section determination, making the result more robust and self-consistent.
@@ IV Sterile Neutrinos
The simplest and most considered explanation of the gallium anomaly is the hypothesis of short-baseline active-sterile neutrino oscillations (see the recent review in Ref. [1]). Figure 5 shows the allowed regions in the space of the effective parameters of 3+1 short-baseline active-sterile oscillations (see Ref. [14]). Our results confirm those obtained in previous works [24, 75, 76, 77, 31, 11, 27, 20], which show that the gallium data, including the recent and most precise BEST data, can be explained by short-baseline active-sterile neutrino oscillations with a large mixing ($0.16\lesssim\sin^{2}2\vartheta_{ee}\lesssim 0.54$ at $2\sigma$, for $\Delta{m}^{2}_{41}\gtrsim 0.98\,\text{eV}^{2}$). Such large mixing is in tension with the bounds obtained from the analysis of reactor antineutrino data [78, 11, 79, 80, 81], the solar neutrino bound [82, 73], and the KATRIN neutrino mass measurement [74]. The solar and KATRIN bounds are shown in Fig. 5 for comparison. The reactor bound is more complicated, because it depends on the reactor antineutrino flux model and on the choice of the data set [78, 11]. These tensions disfavour the sterile neutrino explanation of the gallium anomaly, which is however still considered because so far nobody has found a plausible alternative explanation. [^9] Exotic explanations have been proposed in Refs. [83, 84, 85, 23, 30]. Some have been shown to be in tension with the reactor rate data [20].
@@ V Conclusions
In this work, we critically reassess the status of the gallium anomaly. By employing exact Dirac-Hartree-Fock-Slater wave functions for both the bound and continuum electron states, a new averaging procedure and a refined statistical analysis, we reevaluate the neutrino capture cross-section on 71Ga, showing that a $5.5\sigma$ discrepancy is now achieved with the (p,n) excited state contribution, and $4.7\sigma$ with the ($^{3}\text{He},^{3}\text{H}$) excited state contribution. Such a result is obtained by using the same software and underlying assumptions when evaluating the electron wave functions in both the Fermi function and electron density at the nucleus for the EC process, which enter as a ratio in the cross section. Thus, we avoid possible systematic shifts, making the result more robust and self-consistent. We also present the updated allowed region in the space of the effective parameters of 3+1 active-sterile neutrino mixing. We confirm the need for the explanation of the gallium data of a large active-sterile mixing which is in tension with the results of reactor antineutrino experiments [78, 11, 79, 80, 81], with the solar neutrino bound [82, 73], and the KATRIN neutrino mass measurement [74]. Hence, the solution of the gallium anomaly remains an open problem which may be solved by a new source experiment with a different detector [32, 86, 87].
Acknowledgements.
The authors gratefully acknowledge J. Ekman, J. Bieroń, P. Jönsson for their valuable insights on GRASP and Mehrdad S. Beni for his helpful clarification regarding the electron density at the nucleus.